OpenStudy (anonymous):

28% of what number is 3 1/2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Let x = the number. 28% of what number is 3 1/2 Then 28% of x = 3 1/2.

OpenStudy (anonymous):

12.50

OpenStudy (skullpatrol):

recall $$\text{28%} = \frac{28}{100}$$

OpenStudy (skullpatrol):

$$\text 3\frac{1}{2}=\frac{7}{2}$$

OpenStudy (skullpatrol):

Correct :)